OpenStudy (anonymous):

A square prism container with a base 5 cm by 5 cm is partially filled with water. You drop a clump of metal that weighs 525g into the container, and the water level rises 2cm. What is the denstiy of the metal? assuming the metal is pure, what is the metal? Desities of metals Aluminum-2.81 Copper-8.97 Gold-19.30 lead-11.30 lithium-.54 nickel-8.89 platinum-21.40 pottasoom-.86 silver-10.50 sodium-.97

OpenStudy (anonymous):

the volume of a square prism is = (width)*(length)*(height) = 5*5*5 = 125 cm^3 if the water rises 2cm after you drop the metal, the new volume would be: = 5*5*(5+2) = 5*5*7 = 175 cm^3 We can calculate the density of any metal with: (density) = (mass) / (volume) Hence, (density) = (525g) / (175 cm^3) (density) = 3 g/cm^3 which of the metal you listed has the closes density to 3 g/cm^3 ??

OpenStudy (anonymous):

There is a slight mistake in the above solution: You want to take the new volume and subtract the old volume to get the volume of the metal: 175-125 = 50 cm^3 and divide 525 by that number: 525/50 = 10.5 g/cm^3

OpenStudy (anonymous):

I knew i was missing something!! haha thanks for the catch Bates